I just don't know where to start and how to complete this problem
Events A and B are independent. Find P(A and B) when P(A) = 65% and P(B) = 24%.
P(A and B) = _
Events A and B are independent. Find P(A and B) when P(A) = 65% and P(B) = 24%.

Found 2 solutions by BlGDAVE, ikleyn:
Answer by BlGDAVE(19)
Answer by ikleyn(52778)

.

Every time, when the events A and B are independent, the probability  P(A and B) is the product  P(A)*P(B) :


    P(A and B) = P(A)*P(B) = 0.65*0.24.


Use your calculator to get the value.